What is a if (4,7) and (a,3) are 5 units apart?

Answer by Cromlix(4381)

Distance formula: Square root of(x2 - x1)^2 + (y2 - y1)^2
5 = Sqrt of (a - 4)^2 + (3 - 7)^2
5 = sqrt of a^2 - 8a + 16 + 16
5 = sqrt of a^2 - 8a + 32 = 0
(5)^2 = a^2 - 8a + 32
a^2 - 8a + 32 - 25 = 0
a^2 - 8a + 7 = 0
(a - 7)(a - 1)= 0
a = 7 or a = 1
Hope this helps.
:-)